one number exceeds another by 24.The sum of the numbers is 58.What are the numbers?
Answer by stanbon(57347)

One number exceeds another by 24.
The sum of the numbers is 58.
What are the numbers?
Equations:
x - y = 24
x + y = 58
------
Add and solve for "x":
2x = 82
x = 41
-----------
Solve for "y":
x+y = 58
41 + y = 58
y = 17
====================
Cheers,
Stan H.